OpenStudy (anonymous):

On a grid, the distance between the points located at (3, 2) and (9, 3) is roughly equal to _____.

OpenStudy (anonymous):

you can use the distance formula between two points... \[d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\]

OpenStudy (anonymous):

i got 7

OpenStudy (aum):

\(d=\sqrt{(9-3)^2+(3-2)^2}\)

OpenStudy (anonymous):

it should be...\[\sqrt{37}\]

OpenStudy (anonymous):

whats that like 6

OpenStudy (aum):

\(\sqrt{37} = 6.0827625 \approx 6\)